# Interpreting the Spectro-Temporal Properties of the Black Hole Candidate Swift J151857.0-572147 during its First Outburst in 2024

### v1

@@ Abstract
For the first time, in March 2024, the transient Galactic black hole candidate Swift J151857.0-572147 experienced an outburst. Using publicly available archived Insight-HXMT data, we analyze the timing and spectral features of this source. Through model fitting of the power density spectrum, we were able to extract the properties of quasi-periodic oscillations, and based on those properties, we have determined that the QPOs are of type C. We also conclude that the shock instabilities in the transonic advective accretion processes surrounding black holes may be the source of the QPOs. This shock instability could produce variabilities of flux up to 48 keV, as we checked from the QPO energy dependence. High-frequency QPO is not observed during this period. In the broad energy band of $2-100$ keV, simultaneous data from the three on-board instruments of Insight-HXMT were used to perform the spectral analysis. A combination of models, including broken power-law, multi-color disk-blackbody continuum, interstellar absorption, and reflection in both neutral and ionized medium were needed for spectral fitting to obtain the best fit. We discovered that at the beginning of the analysis period, the source was in an intermediate state and was transitioning toward the softer states based on the spectral features. It has a hydrogen column density of $(4.3-6.9)\times 10^{22}$ cm-2.
X-rays: binary stars; black holes; Stellar accretion disks; Shocks; Compact radiation sources
@@ 1 Introduction
X-ray binaries (XRBs) are quite common and important astronomical binary systems. Since accretion serves as the power source in these systems, it is crucial to understand them (Frank, King & Raine 2002). The primary object in XRBs is a compact object, and the secondary object is a companion star. The compact object may be a neutron star or a black hole, which are both the remnants of stellar bodies. There are multiple categories for XRBs. They fall into two main categories based on the companion’s mass: low-mass X-ray binaries (LMXRBs) and high-mass X-ray binaries (HMXRBs) (Remillard & McClintock 2006). Transient and persistent sources are the other categories into which XRBs are divided, based on the type of variability in their outbursts. While transient sources occasionally exceed detection levels and primarily remain in quiescence, or the dormant state ($L<10^{32}$  erg/s; Hannikainen et al. 2005), flux or counts of persistent sources remain higher than just the detection level most of the time ($L>10^{36}$ erg/s; Chen et al. 1997). These transients experience outbursts that can endure for several weeks or even months (Tetarenko et al. 2016). Though the population of transient HMXBs is increasing, the majority of reported transients are LMXBs (McClintock et al. 2013; Remillard & McClintock 2006 for a review as well). As per Debnath et al. (2010), there are two main categories of BH outbursts based on their nature: slow rise slow decay (SRSD) and fast rise slow decay (FRSD). Zhang et al. (2019) divided outbursts into many types, such as glitter, reflare, multipeak, mini-outburst, or new-outburst, based on their rebrightening characteristics.
The soft multi-color thermal black body and the hard non-thermal power-law components combine to form the spectrum of a black hole. The hard component can be explained by physical scenario in the Comptonizing region, also referred to as the ‘Compton Cloud’, which is the repository of hot electrons (Sunyaev & Titarchuk 1980, 1985). The soft component is modeled as the radiation of the standard Keplerian disk (Shakura & Sunyaev 1973, hereafter SS73). Over time, numerous models have been proposed to describe the hard component of the composite spectrum of a stellar-mass black hole, e.g., the thick disk model (Paczynski & Witta 1980). Even though they did a good job of explaining it, they all made some very specific assumptions. In 1995, Chakrabarti and his collaborators developed the two-component advective flow (TCAF, Chakrabarti & Titarchuk 1995) system, which provided a more comprehensive solution.
A transient black hole experiences many phases and spectral types during an outburst (Remillard & McClintock 2006). Generally speaking, during an outburst, we witness four distinct BH spectral states, which are hard state (HS), hard intermediate state (HIMS), soft intermediate state (SIMS), and soft state (SS), respectively. An outburst usually starts in the HS. After that, it moves to the HIMS and the SIMS as brightness increases. At last, when the source’s brightness peaks, it shifts to the SS. This is referred to as the outburst’s rising phase. The source then transits back to the HS in the opposite cycle as the brightness gradually drops to its minimum. This is referred to as the outburst’s declining phase. To put it briefly, a BH’s spectral state transition forms a hysteresis loop in the order listed below: HS (rising) $\rightarrow$ HIMS (rising) $\rightarrow$ SIMS (rising) $\rightarrow$ SS $\rightarrow$ SIMS (declining) $\rightarrow$ HIMS (declining) $\rightarrow$ HS (declining). A source that experiences all four of the aforementioned often observed spectral states during some outbursts is known as type-I outbursts. In type-II or failed outburst, the source does not go to soft state (Tetarenko et al. 2016).
@@
Understanding temporal aspects is just as critical to comprehending the dynamics of the accreting flow around the BHs as rich spectrum features and variabilities. It has been noted that the light curves exhibit extremely tiny timescale variabilities during an outburst, particularly in the high-energy bands. Variabilities like broadband noise and narrow characteristics in the power density spectrum, or PDS, are imprinted by the Fourier transformation of the light curve (van der Klis 1989). A power-law function is used to describe the broadband noise, which is dispersed over a wide frequency range. Lorentzian profiles can be used to describe the quasi-periodic oscillation (QPO), which is a power peak in restricted frequency ranges. Because of their geometrical origin, low-frequency QPOs (LFQPOs) are frequently detected in BHXRBs. Types A, B, and C are the three categories into which LFQPOs are divided, based on characteristics such as frequency ($\nu$), $Q$-value ($=\nu/\delta\nu$, where $\delta\nu$ is the full width at half maximum, or FWHM), (%) RMS, etc. (Casella et al. 2005). High-frequency QPOs can also be seen in BHXRBs, although it is quite rear. Shock oscillation model (Molteni et al. 1996; Chakrabarti et al. 2005, 2008, 2015), magneto-acoustic waves (Titarchuk et al. 1998), accretion-ejection instability (Tagger & Pellat 1999), Lense-Thirring precession (Stella et al. 1999; Ingram et al. 2009), and other theories have been proposed, though their origin is still up for debate. Shock oscillation model seems to be a more complete one since it can simultaneously describe the temporal and spectral features.
@@
It is thought that spectral and timing properties are related since they originate from the same system and because variations in spectral states can also affect the kinds of QPOs. When examining the two solely in terms of the features of the light curve, such as the hardness ratio, or HR, and the hardness intensity diagram, or HID, a strong association is seen (Homan et al. 2001). Accretion rate ratio intensity diagrams, or ARRIDs, can be used to understand how variations in viscosity or different mass accretion rates interact, as some model-fitted approaches also demonstrated (Chatterjee et al. 2020). Links between spectral and temporal features from pure observational ground can also be established using the RMS-intensity diagram, or RID (Munoz-Darias et al., 2011), and the hardness ratio-intensity diagram, or HRD (Belloni et al., 2005).
@@
First identified by Swift/XRT as a GRB (GRB 20240303A; Kennea et al. 2024), the new Galactic transient Swift J151857.0-572147 was found in Swift Trigger 1218452 (GCN 35849). But thereafter, it was determined to be a Galactic transient due to its constant brightness and location in the Galactic plane. The RA and Dec of the source were determined to be RA(J2000) = $15h18m57.00s$ and Dec(J2000) = $-57d21^{{}^{\prime}}47.9^{{}^{\prime\prime}}$ based on the optimal source localization utilizing XRT instantaneous on-board localization (Kennea et al. 2024). On March 4, 2024, during 15 minutes, from 02:13:13.3 to 02:28:08.9 (MJD 60373.1), follow-up radio observations were conducted using the MeerKAT telescope at $1.28GHz$ (L-band) with a bandwidth of $856$ MHz at flux density of $10$ mJy (Carotenuto et al., 2024; Cowie et al., 2024). The source’s nature was identified as consistent with an X-ray binary in the hard state by using the inverted radio spectrum ($f(\nu)\propto\nu^{\alpha}$, where $\nu\sim+0.5$) in conjunction with the photon index. This suggested that the source might be a black hole or a neutron star. On March 9, 2024, from UT 10:35:10 to UT 11:06:20 (MJD 60378.45), the Australia Telescope Compact Array (ATCA) simultaneously recorded radio observations at frequencies of $5.5$ and $9$ GHz (Saikia et al., 2024). Additionally, their investigation confirmed the source to be a Galactic black hole. Target of opportunity (ToO) was carried out on this source with an exposure of 1000s by Swift/XRT following the ATCA. According to Del Santo et al. (2024), it was discovered that the combination of the phenomenological disk black body (diskbb) and power-law (po) model models describes the spectrum quite well. These discoveries also confirmed that the source is a black hole. The source was detected by INTEGRAL serendipitously on March 8, 9, 10, and 11 of 2024 (Sguera 2024). The 60cm Robotic Eye Mount (REM) telescope observed the source in both optical and near-infrared wavelengths as part of the monitoring program of GRBs (Baglio et al. 2024). Optical measurements of the source were also carried out by the Las Cumbers Observatory (LCO) network (Saikia et al. 2024).
@@
From their Swift/XRT spectral modeling, Kennea et al. (2024) found a column density of $N_{H}=5.6\pm 0.06\times 10^{22}$  cm-2. Additionally, they observed a power-law photon index of $\Gamma=1.78\pm 0.02$. While Burridge et al. (2024) reported that the source’s distance was $4.48^{+0.67}_{-0.47}$ kpc, with an HI absorption towards it, the absence of positive velocity absorption lines towards other sources in the field of the HI absorption for this source puts an upper limit on the distance as $15.64^{+0.77}_{-0.60}$ kpc. The mass and spin parameters of the source are reported to be $\sim 9.2\pm 1.6-10.5\pm 1.8M_{\odot}$ and $\sim 0.65$, while the possible inclination is $\sim 38^{\circ}\pm 8^{\circ}-47^{\circ}\pm 15^{\circ}$ and $0.65$ (Mondal et al. 2024).
@@ 2 Observation and Data Reduction and Analysis
This source has recently been observed by Swift satellite. It is currently being monitored at the time of this manuscript writing by various other X-ray satellites, e.g., NICER, NuSTAR, IXPE, etc. We use X-ray data from China’s first X-ray mission Insight-HXMT (Zhang et al. 2020). After the onset of the outburst, 7 observation IDs were available publicly when we started our analysis. We list the data in Table 1 below.
@@ 2.2 Data Analysis
We conduct spectral and temporal research on the black hole candidate (BHC) Swift J151857.0-572147’s very first outburst in 2024. First, we created $0.01~{}s$ time-binned light curves using data from the LE, ME, and HE modules of the HXMT. The fast Fourier transformation (FFT) in the powspec task of the XRONOS package in the HEASoft software was used to construct the power density spectrum (PDS) based on those light curves. The data from each observation was split up into many intervals, with 8192 new bins in each interval. To create the final PDS, the PDS for each interval must first be generated and then averaged. The PDS is normalized using the Leahy normalization (Leahy et al. 1983). A geometrical rebinning of -1.02 is applied. We used these procedures to look for LFQPOs. Utilizing a Lorentzian model in powspec, we fit the QPO shape and derive QPO properties such as frequency ($\nu_{qpo}$), full-width at half maximum (FWHM) and normalization. We have detected the presence of harmonics in several observations. Additionally, we have derived their properties by Lorentzian model fitting. We fitted the light curves of all the exposures for three energy bands LE ($2-10$ keV), ME ($10-35$ keV), and HE ($27-250$ keV) of the listed (Table 1) 7 observations. We report them next in the result section.
@@
Using these fitting estimations ($\nu_{qpo}$, FWHM, norm), we also estimated some properties of the QPOs that help designate their nature. This will be discussed in the result section.
For HFQPOs, we followed the same procedure, except only using the $1~{}ms$ time-binned light curves for all the exposures for all three energy bands.
We also used all three modules (LE, ME, and HE) for spectral analysis, fitting the broadband data in the $2-150$ keV energy range. First, we tried to do the spectral analysis using a combination of simple disk blackbody and power-law models. However, we did not find an acceptable fit. This is discussed later. The combinations of disk blackbody, broken power-law models yielded the best fit for the data, according to our search. We have employed the tbabs model for interstellar absorption. Since we are simultaneously fitting all three modules, we have included a constant to normalize the three resultant fittings. The following is our best model fit combination: constant*tbabs*(diskbb + broken power-law). We take this as our Model-1. We also tried to analyze spectral data using reflection model pexrav. For that, our model combination reads as: constant*tbabs*(diskbb + pexrav). We take this as our Model-2. We also tried to use the reflection model pexriv which accounts for ionized medium. Thus, constant*tbabs*(diskbb + pexriv) reads as our Model-3. Systematic errors were added to perform spectral analysis, as suggested in the HXMT manual. Not every exposure ID of the specified observation IDs was subjected to spectral analysis. We did not spectrally analyze every exposure, compared to the time analysis. Table 2 indicates the spectrally analyzed exposures with a ‘*’. This is because: in the case of timing analysis, we observed variations in timing properties in a single day, but in the case of spectrum analysis, the properties do not significantly change over a short period. We include them in the section on results. We also tried to spectrally fit the exposures using a relativistic reflection model relxill (Dausar et al. 2016). However, we did not find an acceptable fit for all the observations. This will be discussed later.

### v2

@@ Abstract
The transient Galactic black hole candidate Swift J151857.0-572147 went through an outburst in March 2024 for the first time. Using publicly archived Insight-HXMT data, we have analyzed the timing and spectral properties of the source. We have extracted the properties of the quasi-periodic oscillations (QPOs) by fitting the power density spectrum, which inferred that the QPOs are of type C. We have detected QPOs up to $\sim 48$ keV using an energy dependence study of the QPOs. High-frequency QPO is not observed during this period. We also conclude that the oscillations of the shock in transonic advective accretion flows may be the possible reason for the origin of the QPOs. In the broad energy band of $2-100$ keV, simultaneous data from the three onboard instruments of Insight-HXMT were used to perform spectral analysis. Different combinations of models, including broken power-law, multi-color disk blackbody, interstellar absorption, non-relativistic reflection in both neutral and ionized medium, and relativistic reflection were used to understand the spectral properties during the outburst. We discovered that at the beginning of the analysis period, the source was in an intermediate state and later transitioning toward the soft state based on the spectral parameters. It has a high hydrogen column density, which could be due to some local absorption to the source.
X-rays: binary stars (1811); black holes (162); Stellar accretion disks (1579); Shocks (2086); Compact radiation sources (289)
@@ 1 Introduction
Black hole X-ray binaries (BHXRBs) are quite common and important astronomical binary systems. Since accretion serves as the power source in these systems, it is crucial to understand them (Frank, King & Raine 2002). In BHXRBs, there is a black hole (BH) at the center and a companion star orbiting around. They are classified into two main types based on the companion’s mass: low-mass BHXRBs (LMBHXRBs) and high-mass BHXRBs (HMBHXRBs) (Remillard & McClintock 2006). In LMBHXRBs, the compact object is more massive and known as the primary whereas, the companion is the secondary object. Transient and persistent sources are the other categories into which BHXRBs are divided, based on the type of variability in their outbursts. While BH transient sources occasionally exceed detection levels and primarily remain in quiescence, or the dormant state ($L<10^{32}$  erg/s; Hannikainen et al. 2005 and references therein), the flux or count rate of persistent sources remain higher than the detection level most of the time ($L>10^{36}$ erg/s; Chen et al. 1997 and references therein). Transient sources experience outbursts that can endure for several weeks or even months (Tetarenko et al. 2016). Though the population of transient HMXBs is increasing, the majority of reported transients are LMXBs (McClintock et al. 2013; Remillard & McClintock 2006 for a review as well). Debnath et al. (2010) classified the BH outbursts into two main categories based on their nature: slow rise slow decay (SRSD) and fast rise slow decay (FRSD). Zhang et al. (2019) divided outbursts into several types, such as glitter, reflare, multipeak, mini-outburst, or new-outburst, based on their rebrightening characteristics.
The multicolor thermal black body and the hard non-thermal power-law components combine to form the spectrum of a BH. The origin of the hard component can be explained by the Comptonizing region, also known as the ‘Compton Cloud’, which is the repository of hot electrons (Thorne & Price 1975; Sunyaev & Titarchuk 1980, 1985). Over the years, many models have been put forward to explain the composite spectrum of BHXRBs. For example, the standard disk model (Shakura & Sunyaev 1973, or SS73), the thick disk model (Paczynski & Witta 1980), the advection dominated accretion flow or ADAF (Ichimaru 1977, Narayan & Yi 1994), two-component advective flow or TCAF model (Chakrabarti & Titarchuk 1995), etc. to explain variabilities in BHXRBs. The soft component is modeled as the radiation coming from the disk, which was explained well by the SS73 disk. However, the model could not describe the harder part of the spectrum. The thick disk model could explain the hard power-law part of the spectrum. This model is relevant for high luminosity states, where radiation pressure dominates. However, advection was not taken into consideration in this model. In the ADAF model, the energy generated due to viscous dissipation is advected to the BH and is radiatively inefficient. This model produces a power-law spectrum, however, the direct spectral fitting and explaining the spectral and timing properties are still lacking. Alternatively, the TCAF model solves flow equations and couples them with the radiative transfer processes to explain both spectral and timing properties simultaneously (Chakrabarti & Titarchuk 1995, Molteni et al. 1996, Mondal & Chakrabarti 2013, Chakrabarti et al. 2015).
During the rising phase of an outburst, BH binaries show four distinct BH spectral states starting from hard state (HS) to hard intermediate state (HIMS), soft intermediate state (SIMS), and soft state (SS) (see Remillard & McClintock 2006, for a review). An opposite sequence follows in the declining phase. Briefly, a BH’s spectral state transition forms a hysteresis loop from HS to SS (in the rising phase) and SS to HS (in the declining phase). There is no significant difference in the energy spectrum between HIMS and SIMS. SIMS is classified mostly based on the presence of type-B QPOs. A source that experiences all four of the aforementioned spectral states during an outburst, is known to have a complete outburst. If the source does not go to a soft state, it is known to have a failed outburst (Tetarenko et al. 2016).
@@
Understanding temporal aspects is crucial to understanding the dynamics of the accreting flow around the BHs. The light curves exhibit variabilities in short timescales, in the order of milliseconds to seconds, during an outburst, particularly in the high-energy bands. Variabilities like broadband noise and narrow characteristics in the PDS, can be observed by the Fourier transformation of the light curve (van der Klis 1989; Belloni et al. 2002). A power-law function is used to describe the broadband noise, which is dispersed over a wide frequency range. The broadband noise can also be fitted with a set of zero-centered Lorentzians (Belloni et al. 2002). Lorentzian profiles can be used to describe the QPOs, which is a power peak in the restricted frequency range (Psaltis et al. 1999; Nowak 2000; Belloni et al. 2002). Because of their geometrical origin, low-frequency QPOs (LFQPOs) are frequently detected in BHXRBs. Types A, B, and C are the three categories into which LFQPOs are divided, based on some characteristics such as frequency ($\nu$), $Q$-value ($=\nu/\delta\nu$, where $\delta\nu$ is the full width at half maximum or FWHM), (%) RMS, etc. (Casella et al. 2005). High-frequency QPOs can also be seen in BHXRBs, although it is quite rare. Several models have been put forward to understand the origin of QPOs, such as the shock oscillation model (Molteni et al. 1996; Chakrabarti et al. 2005, 2008, 2015), magneto-acoustic waves (Titarchuk et al. 1998), accretion-ejection instability (Tagger & Pellat 1999), Lense-Thirring precession (Stella et al. 1999; Ingram et al. 2009), precessing inner flow model (Ingram et al. 2009), corrugation modes (Kato & Fukue 1980; Wagoner 1999; Kato 2001; Tsang & Butsky 2013), pressure or accretion rate modes (Cabanac et al. 2010), variable Comptonization or vKompth model (Karpouzas et al. 2021; Bellavita et al. 2022; Garcia et al. 2022; and references therein), outflow model (Reig et al. 2003; Giannios et al. 2004; Kylafis et al. 2020; and references therein), JED-SAD model (Ferreira et al. 1997, 2022; Petruci et al. 2008; Marcel et al. 2019 and references therein), reltrans model (Ingram et al. 2019; Mastroserio et al. 2021; and references therein). However, to date, the origin of QPOs is still a topic of debate.
One of the characteristics of the TCAF model (Chakrabarti & Titarchuk 1995) is the oscillation of the shock. According to the TCAF model, matter supplied by the companion star can have a Keplerian and a sub-Keplerian distribution of angular momentum. The Keplerian component creates a geometrically thin and optically thick accretion disk on the equatorial plane and flows in on a viscous timescale because of its high viscosity. As the critical viscosity of this matter increases, the accretion disk moves inward. The sub-Keplerian one falls radially in on a free-fall timeline and has a less viscous accretion flow. This matter resides both above and below the Keplerian disk. The optically thin sub-Keplerian component produces a shock front at the location where both the centrifugal and gravitational forces balance each other. The shock front is the boundary layer, also known as the CENtrifugal pressure-supported BOundary Layer (CENBOL), which behaves as a so-called Compton cloud. The Keplerian component can explain the soft multicolor blackbody component. The CENBOL region up-scatters the soft photons from the disk and makes them as hard power-law photons. The shock forms farther away from the BH at the beginning of the outburst and the Keplerian disk starts forming, therefore HS is observed. The blackbody photons start increasing as the Keplerian disk moves inward over time. Thereby, cooling increases as a great number of soft photons intercept the CENBOL. Throughout the process, more photons are released, the flux rises, and the source becomes softer. The CENBOL is completely quenched in a high soft state, and the disk approaches the inner most stable orbit. The spectrum becomes soft as a result of only the disk photons now contributing to the radiation. Such profiles of spectral state evolution have been observed in several sources in the TCAF scenario (Mondal et al. 2014, Debnath et al. 2015, Jana et al. 2016, Chatterjee et al. 2020, 2021, 2023).
Furthermore, this model explains the QPO properties in addition to the spectral features and their evolutions. When the infall and cooling timescales become comparable, a bigger CENBOL ($\sim$ a few 100 $r_{S}$, where $r_{S}$ is the Schwartzschild radius) in the HS may result in LFQPOs after satisfying the resonance condition of the oscillation of the shock. The CENBOL slowly shrinks in size as the outburst progresses, and the QPO’s frequency increases (Molteni et al. 1996, Chakrabarti et al. 2015, Garain et al. 2014). Cooling takes over the heating timescale as the SS is approached, the CENBOL gets smaller and eventually quenched, and there is no oscillation. As a result, there is no QPO in the SS (see Mondal et al. 2014, Debnath et al. 2015, Chakrabarti et al. 2015). During the whole evolution path, for different sizes of the CENBOL, different types of QPOs are observed.
@@
While examining the correspondence between the spectral and timing properties solely in terms of the features of the light curve, such as the hardness ratio (HR), and the hardness intensity diagram (HID), a strong association is seen (Homan et al. 2001; Fender et al. 2004; Motta et al. 2011). Additionally, accretion rate ratio intensity diagrams (ARRIDs), can also be used to understand the complete cycle of an outburst from the fundamental accretion flow parameters (Mondal et al. 2014, Jana et al. 2016, Chatterjee et al. 2020). The interlinks between spectral and temporal features from a purely observational ground can also be addressed using the RMS-intensity diagram, or RID (Munoz-Darias et al., 2011) and HRD (Belloni et al., 2005).
@@
First identified by Swift/XRT as a GRB (GRB 20240303A; Kennea et al. 2024), the new Galactic transient Swift J151857.0-572147 was found in Swift Trigger 1218452 (GCN 35849). But thereafter, it was determined to be a Galactic transient due to its constant brightness and location in the Galactic plane. The RA and Dec of the source were determined to be RA(J2000) = $15h18m57.00s$ and Dec(J2000) = $-57d21^{{}^{\prime}}47.9^{{}^{\prime\prime}}$ based on the optimal source localization utilizing XRT instantaneous on-board localization (Kennea et al. 2024). On March 4, 2024, during 15 minutes, from 02:13:13.3 to 02:28:08.9 (MJD 60373.1), follow-up radio observations were conducted using the MeerKAT telescope at $1.28GHz$ (L-band) with a bandwidth of $856$ MHz at a flux density of $10$ mJy (Carotenuto et al., 2024; Cowie et al., 2024). The source’s nature was identified as consistent with an X-ray binary in the hard state by using the inverted radio spectrum ($f(\nu)\propto\nu^{\alpha}$, where $\nu\sim+0.5$) in conjunction with the photon index. This suggested that the source might be a black hole or a neutron star. On March 9, 2024, from UT 10:35:10 to UT 11:06:20 (MJD 60378.45), the Australia Telescope Compact Array (ATCA) simultaneously recorded radio observations at frequencies of $5.5$ and $9$ GHz (Saikia et al., 2024). Additionally, their investigation confirmed the source to be a Galactic black hole. Target of Opportunity (ToO) was carried out on this source with an exposure of 1000s by Swift/XRT following the ATCA, as reported by Del Santo et al. (2024). According to Del Santo et al. (2024), it was discovered that the combination of the phenomenological disk black body (diskbb) and power-law (po) models describes the spectrum quite well. These discoveries also confirmed that the source is a black hole. The source was detected by INTEGRAL serendipitously on March 8, 9, 10, and 11 of 2024 (Sguera 2024). The 60cm Robotic Eye Mount (REM) telescope observed the source in both optical and near-infrared wavelengths as part of the monitoring program of GRBs (Baglio et al. 2024). Optical measurements of the source were also carried out by the Las Cumbers Observatory (LCO) network (Saikia et al. 2024).
@@
From the Swift/XRT spectral modeling, Kennea et al. (2024) found a column density of $N_{H}=5.6\pm 0.06\times 10^{22}$  cm-2. Additionally, they observed a power-law photon index of $\Gamma=1.78\pm 0.02$. While Burridge et al. (2024) reported that the source’s distance was $4.48^{+0.67}_{-0.47}$ kpc, with an HI absorption towards it, the absence of positive velocity absorption lines towards other sources in the field of the HI absorption for this source puts an upper limit on the distance as $15.64^{+0.77}_{-0.60}$ kpc. Mondal et al. (2024) reported the mass of the source to be $\sim 9.2\pm 1.6-10.5\pm 1.8M_{\odot}$, estimated using the JeTCAF (Jet in TCAF) model (Mondal & Chakrabarti 2021), where the mass of the BH is a parameter and the distance to the source is a scale factor taken care of by the model normalization. Authors also estimated the possible disk inclination of $\sim 35^{\circ}\pm 7^{\circ}-46^{\circ}\pm 15^{\circ}$ with an average spin parameter of $0.65$ estimated using the KERRBB model (Li et al. 2005) for a given distance of 10 kpc. Since there is no confirmed distance estimation for this source, the authors have adopted an average value from the broad range of proposed distances in the literature. Therefore, the estimation of the spin may vary for different source distance values, requiring further modeling with a confirmed source distance.
@@ 2 Observation and Data Reduction and Analysis
This source has recently been observed by Swift satellite and confirmed to be a black hole candidate on March 10, 2024 (Del Santo et al. 2024). After its discovery and confirmation, it was monitored by various other X-ray satellites, for example, NICER, NuSTAR, IXPE, etc. We use X-ray data from China’s first X-ray satellite mission Insight-HXMT (Zhang et al. 2020). After the onset of the outburst, 7 observation IDs were available publicly when we started our analysis. We list the data in Table 1 below.
@@ 2.2 Data Analysis
We conducted spectral and temporal research on the black hole candidate (BHC) Swift J151857.0-572147’s very first outburst in 2024. First, we created $0.01~{}s$ time-binned light curves using data from the LE, ME, and HE modules of the HXMT. The fast Fourier transformation (FFT) in the powspec task of the XRONOS package in the HEASoft software was used to construct the power density spectrum (PDS) based on those light curves. The data from each observation was split up into many intervals, with 8192 new bins in each interval. To create the final PDS, the PDS for each interval must first be generated and then averaged. The PDS is normalized using the Leahy normalization (Leahy et al. 1983). A geometrical rebinning of -1.02 is applied. We used these procedures to look for LFQPOs. Initially, the analysis was done without subtracting the white noise. With the concern that it may affect the QPO properties, we have rechecked the analysis by considering the white noise subtraction by running the powspec task to produce the PDS using $norm=-1$. In both the cases, using the combination of either power-law, constant, and Lorentzian or multiple Lorentzian models, we fit all the features of the full PDS continuum from 0.01 to 50 Hz, including the fundamental QPO and harmonic (if present) features. We also estimated several QPO properties such as frequency ($\nu_{qpo}$), full-width at half maximum (FWHM), and normalization. We found that the normalization of the fundamental QPO barely changes while we use white noise subtraction to that when we don’t consider white noise subtraction. The normalization value stays well within the error range. The effect of white noise is negligible due to the fact that the signal-to-noise ratio of the narrow fundamental QPO feature is high enough to make the contribution from the white noise negligible. We report the results in the next section. We found a sharp peak at the position of $2\times\nu_{qpo}$ Hz for some observations, which happen to be the harmonic of the fundamental QPO. We used an additional Lorentzian model to fit the harmonic peak. We have extracted their properties from the Lorentzian model fitting. We fitted the PDS of all the exposures for three energy bands LE ($2-10$ keV), ME ($10-35$ keV), and HE ($27-250$ keV) of the listed 7 observations (Table 1). We discuss them later in the result section.
@@
Using these QPO properties ($\nu_{qpo}$, FWHM, norm), we also estimated some properties of the QPOs, like Q-value, RMS (%), that help designate their nature. We have also estimated the QPO significance in the result section following Sreehari et al. (2019).
We also used all three modules (LE, ME, and HE) for spectral analysis, fitting the broadband data in the $2-100$ keV energy range. First, we tried to do the spectral analysis using a combination of simple disk blackbody and power-law models. However, we did not find an acceptable fit. This is discussed in the Appendix. The combinations of disk blackbody, broken power-law models yielded the best fit for the data, according to our search. We have employed the tbabs model for interstellar absorption. Since we are simultaneously fitting all three modules, we have included a constant to normalize the three resultant fittings. For LE instrument, we have frozen the constant value to 1 and let the other two constants vary for ME and HE instruments. These are given in each table of the spectral fitting results in the Appendix section, where constant1 and constant2 are for the ME and HE instruments, respectively. The following is our best model fit combination: constant*tbabs*(diskbb + broken power-law). We take this as our Model-1. We also tried to analyze spectral data using reflection model pexrav. For that, our model combination reads as: constant*tbabs*(diskbb + pexrav). We take this as our Model-2. We also tried to use the reflection model pexriv which accounts for ionized medium. Thus, constant*tbabs*(diskbb + pexriv) reads as our Model-3. We also performed the spectral fitting using the relativistic relxill model (Dauser et al. 2016), using the model combination as constant*tbabs(diskbb + relxill). We take this as Model-4. Systematic errors were added to perform spectral analysis, as suggested in the HXMT manual. Not every exposure ID of the specified observation IDs was subjected to spectral analysis. We did not spectrally analyze every exposure, compared to the time analysis. Table 2 indicates the spectrally analyzed exposures with a ‘*’. This is because: in the case of timing analysis, we observed variations in timing properties in a single day, but in the case of spectrum analysis, the properties do not significantly change over a short period. We include them in the section on results.